0.790 g sample of an unknown monoprotic acid is dissolved in water and titrated with standardized barium hydroxide. The equivalence point in the titration is reached after the addition of 27.3 mL of 0.261 M barium hydroxide to the sample of the unknown acid. Calculate the molar mass of the acid.

Give your answer to three significant figures. 


1
Expert's answer
2020-10-26T14:56:57-0400

N1V1=N2V2=n2,

N1=N(Ba(OH)2)=2C(Ba(OH)2)=2×0.261=0.522(mol-eq/L),

n2=0.522×0.0273=0.0143(mol),

M=m/n=0.790/0.0143=55.2448(g/mol)